# Scalable Valuation of Human Feedback through Provably Robust Model Alignment

### v1

@@ E.4 Proof for Corollary 1
According to the fact that $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\sigma(-g_{\theta^{*}}(s_{\textrm{flip}}))=1$ and $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\sigma(g_{\theta^{*}}(s_{\textrm{flip}}))^{\gamma}=0$ for $0<\gamma<\infty$, we can see that $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}}^{(\gamma)}(s_{\textrm{flip}})]=\mathbb{E}_{p(s_{\textrm{flip}})}[\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}F_{\theta^{*}}^{(\gamma)}(s_{\textrm{flip}})]0$ since $\nabla_{\theta}\mathcal{L}(s_{\textrm{flip}},\pi_{\theta^{*}})$ is bounded by $2C$ following the same discussion in Corollary 2. Therefore, whether the robustness property holds or not depends on the behavior of $H_{\theta^{*}}^{(\gamma)}(s)$ and $H_{\theta^{*}}^{(1+\gamma)}(s)$ included in $C^{(\gamma)}_{\theta^{*}}(s)$ of Eq. (21).
Following the discussion in Corollary 2, we can see that $\nabla_{\theta}\mathcal{L}(s,\pi_{\theta^{*}})$ is $(\sigma(-g_{\theta^{*}}(s))\cdot L)$-Lipschits continuous in $\theta$. Since $10<\sigma(g_{\theta^{*}}(s))^{\gamma}<$ 1, the term $F_{\theta^{*}}^{(\gamma)}(s)=\sigma(g_{\theta^{*}}(s))^{\gamma}\nabla_{\theta}\mathcal{L}(s,\pi_{\theta^{*}})$ and $F_{\theta^{*}}^{(1+\gamma)}(s)=\sigma(g_{\theta^{*}}(s))^{1+\gamma}\mathcal{L}(s,\pi_{\theta^{*}})$ are $(\sigma(g_{\theta^{*}}(s))^{\gamma}\sigma(-g_{\theta^{*}}(s))\cdot L)$-Lipschits continuous and $(\sigma(g_{\theta^{*}}(s))^{(1+\gamma)}\sigma(-g_{\theta^{*}}(s))\cdot L)$-Lipschits continuous in $\theta$, respectively. Thus, by using Lemma 1.2.2 in Nesterov [75], we have $0<\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}^{(\gamma)}(s)]\leq\sigma(g_{\theta^{*}}(s))^{\gamma}\sigma(-g_{\theta^{*}}(s))\cdot L<\infty$ and $0<\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}^{(1+\gamma)}(s)]\leq\sigma(g_{\theta^{*}}(s))^{1+\gamma}\sigma(-g_{\theta^{*}}(s))\cdot L<\infty$. This completes the proof. ∎
@@ F.2.2 Extended model for estimating the contamination ratio
$\operatorname*{\mathop{\rm argmin}}_{\theta}\min_{\xi}D_{H}[\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s})]=\operatorname*{\mathop{\rm argmin}}_{\theta}\min_{\xi}S_{\gamma}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s})).$ (22)
@@
$S_{\gamma}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}))$ $=\phi\bigg{(}\frac{\mathbb{E}_{\tilde{p}_{\mathcal{D}}^{(\epsilon)}}[m_{\eta}^{\gamma}(\tilde{s})]}{\int m_{\eta}^{1+\gamma}(\tilde{s})\mathrm{d}\tilde{s}}\bigg{)}\bigg{(}\int m_{\eta}^{1+\gamma}(\tilde{s})\mathrm{d}\tilde{s}\bigg{)}$
$=\phi\bigg{(}\frac{\mathbb{E}_{\tilde{p}_{\mathcal{D}}^{(\epsilon)}}[\sigma(g_{\theta}(\tilde{s}))^{\gamma}]}{\xi\cdot\int\sigma(g_{\theta}(\tilde{s}))^{1+\gamma}\mathrm{d}\tilde{s}}\bigg{)}\bigg{(}\xi^{1+\gamma}\cdot\int\sigma(g_{\theta}(\tilde{s}))^{1+\gamma}\mathrm{d}\tilde{s}\bigg{)}$
$\geq-\bigg{(}\frac{\mathbb{E}_{\tilde{p}_{\mathcal{D}}^{(\epsilon)}}[\sigma(g_{\theta}(\tilde{s}))^{\gamma}]}{\xi\cdot\int\sigma(g_{\theta}(\tilde{s}))^{1+\gamma}\mathrm{d}\tilde{s}}\bigg{)}^{1+\gamma}\bigg{(}\xi^{1+\gamma}\cdot\int\sigma(g_{\theta}(\tilde{s}))^{1+\gamma}\mathrm{d}\tilde{s}\bigg{)}$
@@
$=S_{\mathrm{PS}}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|\sigma(g_{\theta}(\tilde{s})))=-\exp\bigg{\{}-\gamma(1+\gamma)\cdot S_{\log}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|\sigma(g_{\theta}(\tilde{s})))\bigg{\}},$ (23)
where the third inequality comes from the fact that $\phi(h)\geq-h^{1+\gamma}$ for all $h\geq 0$, the term $S_{\mathrm{PS}}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|\sigma(g_{\theta}(\tilde{s})))$ in the forth line is called as the pseudo-spherical (PS) score, and $S_{\log}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|\sigma(g_{\theta}(\tilde{s})))$ is the $\gamma$-score associated with $\gamma$-divergence defined as
@@
From the Eq. (F.2.2), we can see that the lower bound of $S_{\gamma}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}))$ is independent of $\xi$. Therefore, the optimal solution of $\xi^{*}=\min_{\xi}S_{\gamma}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}))$ is obtained as the equality condition of Eq. (F.2.2), that is,
@@
At the end of this section, we verify that $\xi^{*}$ in Eq. (25) serves as an estimate of the contamination rate. When robust model alignment using Hölder-DPO effectively reduce the effects of contamination and the value of $\sigma_{\theta}(g_{\theta}(\tilde{s}))$ closely approximates the target distribution $p_{\mathcal{D}}(s)$, i.e., $\sigma_{\theta}(g_{\theta}(\tilde{s}))\approx p_{\mathcal{D}}(s)$, we have:
@@ F.4 Choice of $\phi$
Under Assumption 1 and Definition 1. Then, under the extended model $m_{\eta}$, the objective $S_{\gamma}(p^{(\epsilon)}_{\widetilde{\mathcal{D}}}\|m_{\eta}(\tilde{s}))$ can be approximated around $\theta=\theta^{*}$ as
@@
$S_{\gamma}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}))$ $=\underbrace{-\frac{\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[m_{\eta}(\tilde{s})^{\gamma}]}{\left(\mathbb{E}_{m_{\eta}}[m_{\eta}(\tilde{s})^{\gamma}]\right)^{\frac{\gamma}{1+\gamma}}}}_{\eqqcolon S_{\mathrm{PS}}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}))}$
@@
Under Assumption 1, the optimal solution of $\operatorname*{\mathop{\rm argmin}}_{\theta}S_{\gamma}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}))$ will be close to that of $\operatorname*{\mathop{\rm argmin}}_{\theta}S_{\mathrm{PS}}(p_{\mathcal{D}}(s)\|\sigma(g_{\theta}(s)))$. This implies that Hölder-DPO with $\phi(h)=-h^{1+\gamma}$ is robust to heavy contamination, since it does not require the contamination ratio $\epsilon$ to be small.
@@
$S_{\mathrm{PS}}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}))$ $=-\frac{\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[m_{\eta}(\tilde{s})^{\gamma}]}{\left(\mathbb{E}_{m_{\eta}}[m_{\eta}(\tilde{s})^{\gamma}]\right)^{\frac{\gamma}{1+\gamma}}}$
@@
$=S_{\mathrm{PS}}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|\sigma(g_{\theta}(\tilde{s}))),$
@@
$\operatorname*{\mathop{\rm argmin}}_{\eta=\{\xi,\theta\}}S_{\mathrm{PS}}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}))=\operatorname*{\mathop{\rm argmin}}_{\theta}S_{\mathrm{PS}}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|\sigma(g_{\theta}(\tilde{s}))),$
@@
$S_{\gamma}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}))$ $=\underbrace{\gamma\mathbb{E}_{m_{\eta}}[m_{\eta}(\tilde{s})^{\gamma}]-(1+\gamma)\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[m_{\eta}(\tilde{s})^{\gamma}]}_{\eqqcolon S_{\mathrm{DP}}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}))}$
@@
If $\mathbb{E}_{p(s_{\textrm{flip}})}[\sigma(g_{\theta}(s_{\textrm{flip}}))^{\gamma}]\approx 0$ around $\theta=\theta^{*}$, then the optimal solution of $\operatorname*{\mathop{\rm argmin}}_{\eta}S_{\gamma}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}))$ is close to that of $\operatorname*{\mathop{\rm argmin}}_{\eta}S_{\mathrm{DP}}((1-\epsilon)p_{\mathcal{D}}\|m_{\eta}(s))$. Recalling that the DP divergence is strictly proper over the set of non-negative functions [52], this implies that minimizing the DP divergence with the extended model allows for estimation of both the target parameter $\theta^{*}$ and the clean-data ratio $1-\epsilon$.
Appendix G IF Analysis for the DPO variants
G.1 cDPO do not satisfy the redescending property
The objective of cDPO [73] is as follows:
~$\widetilde{\mathcal{L}}_{\mathrm{cDPO}}(\pi_{\theta};\pi_{\mathrm{ref}})\coloneqq(1-c)\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(g_{\theta}(\tilde{s}))]-c\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(-g_{\theta}(\tilde{s}))].$ (28)
We first show the IF for the cDPO.
Theorem 6.
$\mathrm{IF}_{\mathrm{cDPO}}(x,\theta,p_{\mathcal{D}})=\bigg{|}\bigg{(}\mathbb{E}_{p_{\mathcal{D}}}\left[H_{\theta^{*}}^{(\textrm{cDPO})}(s)\right]\bigg{)}^{-1}\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}}^{(\textrm{cDPO})}(s_{\textrm{flip}})]\bigg{|},$ (29)
Proof.
The proof follows from the same argument as in Theorem 8, ignoring the $(1-2c)$ term in the denominator. ∎
Corollary 3 (cDPO is not robust).
Suppose that the policy gradient $\nabla_{\theta}\log\pi_{\theta}(y\mid x)$ is bounded by $C$ and satisfies $L$-Lipchitz in $\theta$, where $0<C<\infty$ and $0<L<\infty$. Let $10\leq c<$ 1. Then, under Theorem 6, the IF of cDPO do not satisfy the robustness condition in Definition 2, i.e., $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathrm{IF}_{\mathrm{cDPO}}(x,\theta,p_{\mathcal{D}})\neq 0$.
Proof.
By following the proof in Corollary 5 and ignoring the $(1-2c)$ term in the denominator, we have $0<\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}^{(\textrm{cDPO})}(s)]\leq\xi_{\theta^{*}}(s)\cdot L<\infty$ and thus $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathrm{IF}_{\mathrm{cDPO}}(x,\theta,p_{\mathcal{D}})\leq 2C(1-c)/L^{\prime}$, where $0<L^{\prime}\leq\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}(s)]$ and we use the fact that $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\xi_{\theta^{*}}(s_{\textrm{flip}})=1-c$. The fact $0<2C(1-c)/L^{\prime}<\infty$ according to $0<C<\infty$, $10\leq c<$ 1, and $0<L^{\prime}<\infty$ completes the proof. ∎
G.2 IPO do not satisfy the redescending property
The objective of IPO [36] is as follows:
~$\widetilde{\mathcal{L}}_{\mathrm{IPO}}(\pi_{\theta};\pi_{\mathrm{ref}})\coloneqq\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}\bigg{[}\bigg{(}\sigma\bigg{(}\frac{g_{\theta}(\tilde{s})}{\beta}\bigg{)}-\frac{1}{2\beta}\bigg{)}^{2}\bigg{]}.$ (30)
We first show the IF for the IPO.
Theorem 7.
$\mathrm{IF}_{\mathrm{IPO}}(x,\theta,p_{\mathcal{D}})=\bigg{|}\bigg{(}\mathbb{E}_{p_{\mathcal{D}}}\left[H_{\theta^{*}}^{(\textrm{IPO})}(s)\right]\bigg{)}^{-1}\mathbb{E}_{p(s_{\textrm{flip}})}[F_{\theta^{*}}^{(\textrm{IPO})}(s_{\textrm{flip}})]\bigg{|},$ (31)
where $F_{\theta^{*}}^{(\textrm{IPO})}(s_{\textrm{flip}})\coloneqq\bigg{(}\frac{g_{\theta}(s_{\textrm{flip}})}{\beta}-\frac{1}{2\beta}\bigg{)}\bigg{(}\nabla_{\theta}\log\pi_{\theta^{*}}(y_{\textrm{win}}^{\textrm{flip}}\mid x)-\nabla_{\theta}\log\pi_{\theta^{*}}(y_{\textrm{lose}}^{\textrm{flip}}\mid x)\bigg{)}$.
Proof.
$\nabla_{\theta}\widetilde{\mathcal{L}}_{\mathrm{IPO}}(\pi_{\theta};\pi_{\mathrm{ref}})=\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}\bigg{[}\bigg{(}\frac{g_{\theta}(\tilde{s})}{\beta}-\frac{1}{2\beta}\bigg{)}\bigg{(}\nabla_{\theta}\log\pi_{\theta}(\tilde{y}_{\textrm{win}}\mid\tilde{x})-\nabla_{\theta}\log\pi_{\theta}(\tilde{y}_{\textrm{lose}}\mid\tilde{x})\bigg{)}\bigg{]}.$
Corollary 4 (IPO is not robust).
Suppose that the policy gradient $\nabla_{\theta}\log\pi_{\theta}(y\mid x)$ is bounded by $C$ and satisfies $L$-Lipchitz in $\theta$, where $0<C<\infty$ and $0<L<\infty$. Let $g_{\theta^{*}}(s)$ be bounded over $p_{\mathcal{D}}(s)$. Then, under Theorem 7, the IF of IPO do not satisfy the robustness condition in Definition 2, i.e., $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathrm{IF}_{\mathrm{IPO}}(x,\theta,p_{\mathcal{D}})=\infty$.
Proof.
From the positive definite assumption in Theorem 7 and the $L$-Lipschitsness assumption of $\nabla_{\theta}\log\pi_{\theta}(y\mid x)$ in $\theta$, we can see that $F_{\theta^{*}}^{(\textrm{IPO})}(s)$ is $((|g_{\theta^{*}}(s)|/\beta-1/2\beta)\cdot L)$-Lipschits continuous in $\theta$. We thus have $0<\mathbb{E}_{p_{\mathcal{D}}}[H_{\theta^{*}}^{(\textrm{rDPO})}(s)]\leq(|g_{\theta^{*}}(s)|/\beta-1/2\beta)\cdot L<\infty$ under $0\leq|g_{\theta^{*}}(s)|<\infty$, where we use Lemma 1.2.2 in Nesterov [75]. However, according to the fact that $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}g_{\theta^{*}}(s_{\textrm{flip}})=-\infty$, we can see that $\lim_{r_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathrm{IF}_{\mathrm{IPO}}(x,\theta,p_{\mathcal{D}})=\infty$. This completes the proof. ∎
G.3 rDPO do not satisfy the redescending property

### v2

@@ E.4 Proof for Corollary 1
From Theorem 4, the IF for Hölder-DPO is given by:
$\mathrm{IF}_{\mathrm{H-DPO}}=-\bigg(C^{(\gamma)}_{\theta^{*}}(s)\bigg)^{-1}\cdot\gamma\phi^{\prime}(h_{\theta^{*}})\mathbb{E}_{p(s_{\textrm{flip}})}[F^{(\gamma)}_{\theta^{*}}(s_{\textrm{flip}})].$
First, we analyze the denominator $C^{(\gamma)}_{\theta^{*}}(s)$. Following Theorem 5, we can see that the Hessian-related term $C^{(\gamma)}_{\theta^{*}}(s)$ is positive definite. Let $L^{\prime}=\lambda_{\mathrm{min}}(C^{(\gamma)}_{\theta^{*}}(s))>0$ be its minimum eigenvalue. Then the norm of its inverse is bounded: $\|(C^{(\gamma)}_{\theta^{*}}(s))^{-1}\|\leq 1/L^{\prime}$.
Next, we analyze the numerator, $\mathbb{E}_{p(s_{\textrm{flip}})}[F^{(\gamma)}_{\theta^{*}}(s_{\textrm{flip}})]$, where $F^{(\gamma)}_{\theta^{*}}(s_{\textrm{flip}})\coloneqq\sigma(g_{\theta^{*}}(s_{\textrm{flip}}))^{\gamma}(1-\sigma(g_{\theta^{*}}(s_{\textrm{flip}})))\nabla_{\theta}g_{\theta^{*}}(s_{\textrm{flip}})$. From the assumption that $\|\nabla_{\theta}\log\pi_{\theta}\|\leq C$, the term $\|\nabla_{\theta}g_{\theta^{*}}\|$ is also bounded by $C^{\prime}=2\beta C$.
We now take the limit required by Definition 2:
$\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}$ $\|\mathrm{IF}_{\mathrm{H-DPO}}\|$
$\leq\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\bigg\|\bigg(C^{(\gamma)}_{\theta^{*}}(s)\bigg)^{-1}\cdot\gamma\phi^{\prime}(h_{\theta^{*}})\bigg\|\cdot\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\bigg\|\mathbb{E}_{p(s_{\textrm{flip}})}[F^{(\gamma)}_{\theta^{*}}(s_{\textrm{flip}})]\bigg\|$
$\leq K\cdot\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathbb{E}_{p(s_{\textrm{flip}})}\bigg[\underbrace{\sigma(g_{\theta^{*}}(s_{\textrm{flip}}))^{\gamma}}_{\to 0}\cdot\underbrace{(1-\sigma(g_{\theta^{*}}(s_{\textrm{flip}})))}_{\to 1}\cdot\underbrace{\|\nabla_{\theta}g_{\theta^{*}}\|}_{\leq C^{\prime}}\bigg],$
where $K=(1/L^{\prime})\cdot|\gamma\phi^{\prime}(h_{\theta^{*}})|$ is a finite non-zero constant.
By the bounded convergence theorem, we can move the limit inside the expectation:
$\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\|\mathrm{IF}_{\mathrm{H-DPO}}\|$ $\leq K\cdot\mathbb{E}_{p(s_{\textrm{flip}})}\bigg[\lim_{\hat{r}_{\theta^{*}}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\Big(\sigma(g_{\theta^{*}})^{\gamma}\cdot(1-\sigma(g_{\theta^{*}}))\cdot\|\nabla_{\theta}g_{\theta^{*}}\|\Big)\bigg]$
$\leq K\cdot\mathbb{E}_{p(s_{\textrm{flip}})}\left[0\cdot 1\cdot C^{\prime}\right]$
$=0.$
The IF of Hölder-DPO converges to 0, satisfying the redescending property. This completes the proof. ∎
@@ F.2.2 Extended model for estimating the contamination ratio
$).\displaystyle\operatorname*{\mathop{\rm argmin}}_{\theta}\min_{\xi}D_{H}[\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s})]=\operatorname*{\mathop{\rm argmin}}_{\theta}\min_{\xi}S_{\gamma}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}$)). (22)
@@
[…]
@@
[…]
@@
From the Eq. (F.2.2), we can see that the lower bound of $)S_{\gamma}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}$)) is independent of $\xi$. Therefore, the optimal solution of $)\xi^{*}=\min_{\xi}S_{\gamma}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}$)) is obtained as the equality condition of Eq. (F.2.2), that is,
@@
At the end of this section, we verify that $\xi^{*}$ in Eq. (25) serves as an estimate of the contamination rate. When robust model alignment using Hölder-DPO effectively reduce the effects of contamination and the value of $)\sigma_{\theta}(g_{\theta}(\tilde{s}$)) closely approximates the target distribution $p_{\mathcal{D}}(s)$, i.e., $\sigma_{\theta}(g_{\theta}(\tilde{s}))\approx p_{\mathcal{D}}(s)$, we have:
@@ F.4 Choice of $\phi$
Under Assumption 1 and Definition 1. Then, under the extended model $m_{\eta}$, the objective $)S_{\gamma}(p^{(\epsilon)}_{\widetilde{\mathcal{D}}}\|m_{\eta}(\tilde{s}$)) can be approximated around $\theta=\theta^{*}$ as
@@
$)\displaystyle S_{\gamma}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}$)) $=\underbrace{-\frac{\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[m_{\eta}(\tilde{s})^{\gamma}]}{\left(\mathbb{E}_{m_{\eta}}[m_{\eta}(\tilde{s})^{\gamma}]\right)^{\frac{\gamma}{1+\gamma}}}}_{\eqqcolon S_{\mathrm{PS}}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}))}$
@@
Under Assumption 1, the optimal solution of $)\operatorname*{\mathop{\rm argmin}}_{\theta}S_{\gamma}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}$)) will be close to that of $\operatorname*{\mathop{\rm argmin}}_{\theta}S_{\mathrm{PS}}(p_{\mathcal{D}}(s)\|\sigma(g_{\theta}(s)))$. This implies that Hölder-DPO with $\phi(h)=-h^{1+\gamma}$ is robust to heavy contamination, since it does not require the contamination ratio $\epsilon$ to be small.
@@
$)\displaystyle S_{\mathrm{PS}}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}$)) $=-\frac{\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[m_{\eta}(\tilde{s})^{\gamma}]}{\left(\mathbb{E}_{m_{\eta}}[m_{\eta}(\tilde{s})^{\gamma}]\right)^{\frac{\gamma}{1+\gamma}}}$
@@
[…]
@@
[…]
@@
$)\displaystyle S_{\gamma}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}$)) $=\underbrace{\gamma\mathbb{E}_{m_{\eta}}[m_{\eta}(\tilde{s})^{\gamma}]-(1+\gamma)\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[m_{\eta}(\tilde{s})^{\gamma}]}_{\eqqcolon S_{\mathrm{DP}}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}))}$
@@
If $\mathbb{E}_{p(s_{\textrm{flip}})}[\sigma(g_{\theta}(s_{\textrm{flip}}))^{\gamma}]\approx 0$ around $\theta=\theta^{*}$, then the optimal solution of $)\operatorname*{\mathop{\rm argmin}}_{\eta}S_{\gamma}(\tilde{p}_{\mathcal{D}}^{(\epsilon)}\|m_{\eta}(\tilde{s}$)) is close to that of $\operatorname*{\mathop{\rm argmin}}_{\eta}S_{\mathrm{DP}}((1-\epsilon)p_{\mathcal{D}}\|m_{\eta}(s))$. Recalling that the DP divergence is strictly proper over the set of non-negative functions [53], this implies that minimizing the DP divergence with the extended model allows for estimation of both the target parameter $\theta^{*}$ and the clean-data ratio $1-\epsilon$.
Appendix G IF Analysis for the DPO variants (summarized in Theorem 3)
G.1 rDPO do not satisfy the redescending property